# Cointegration with Occasionally Binding Constraints

### v3

@@ Appendix D Transposition to the structural form
Here we verify the claims made in part (i) of each of Remarks 4.14.3, and complete the proof of 4.4. We suppose that DGP holds, so that $(y_{t},x_{t})$ are generated by the CKSVAR
@@
By 2.1, if $(\tilde{y}_{t},\tilde{x}_{t})$ are constructed from $(y_{t},x_{t})$ via (LABEL:canon-vars), then $(\tilde{y}_{t},\tilde{x}_{t})$ follow a canonical CKSVAR, which we denote here as
@@
We use tildes to distinguish the series and parameters of the canonical form (LABEL:cksvar-structural), from the original CKSVAR (LABEL:cksvar-canon) from which they were derived; the mapping between these is given in (LABEL:canon-vars)–(LABEL:Q-canon).
Our approach here is as follows. We show that if DGP, CVAR, DET (or DET) and a suitably modified form of CO(i) (resp. CO(ii), CO(iii)) hold for $(y_{t},x_{t})$, then DGP, CVAR, DET (or DET) and CO(i) (resp. CO(ii), CO(iii)) also hold for $(\tilde{y}_{t},\tilde{x}_{t})$. Since ERR (or ERR) trivially carry over from $u_{t}$ to $\tilde{u}_{t}$, this permits Theorems 4.14.3 to be applied to $(\tilde{y}_{t},\tilde{x}_{t})$: their implications for $(y_{t},x_{t})$ are then derived by inverting the mapping (LABEL:canon-vars)–(LABEL:Q-canon).
@@
By 2.1, a CKSVAR satisfying DGP has a canonical form, which relates to the original model via (LABEL:canon-vars)–(LABEL:canon-polys). Recall the definition of $(P^{\pm})^{-1}$ given in (LABEL:Ppm). Since
@@
as may be verified using the partitioned inverse formula, (LABEL:canon-polys) may be equivalently stated as
@@
and thus (i) and (ii) hold. Since $\Pi^{\pm}=-\Phi^{\pm}(1)$ and $\tilde{\Pi}^{\pm}=-\tilde{\Phi}^{\pm}(1)$, (iii) follows immediately upon taking $\lambda=1$ in (LABEL:phitilde), and the fact that
@@
(i). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(i), except with CO(i).2 replaced by (ii). We must verify that $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i). $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i).1 by D.1(iii); CO(i).2 since $(y_{t},x_{t})$ satisfies (ii); and CO(i).4 via 2.1.
It remains therefore to verify CO(i).3. We note that as a further consequence of D.1(iii),
@@
Hence, using D.1(iii) again,
@@
It follows that 4.1 applies to $(\tilde{y}_{t},\tilde{x}_{t})$. Regarding the conclusions of that theorem, note that in this case $\tilde{{\cal Y}}_{0}$ and ${\cal Y}_{0}$ must be non-negative. Hence $\tilde{{\cal Z}}_{0}=(P^{+})^{-1}{\cal Z}_{0}$ by (LABEL:canon-vars), and so by 2.1
@@
where $=_{(1)}$ follows by (LABEL:kaptilde), and $=_{(2)}$ by (LABEL:kap1tilde) and $e_{1}^{\mathsf{T}}(P^{+})^{-1}=\bar{\phi}_{0,yy}^{+}e_{1}^{\mathsf{T}}$. It follows immediately that $\beta^{+\mathsf{T}}z_{t}$ and $y_{t}^{-}$ are $I^{\ast}(0)$. Since $\tilde{y}_{t}^{-}\sim I^{\ast}(0)$, it follows that $\Delta\tilde{y}_{t}^{-}\sim I^{\ast}(0)$, and therefore so too is $\Delta\tilde{y}_{t}^{+}=\Delta\tilde{y}_{t}-\Delta\tilde{y}_{t}^{-}$. Hence
@@
(ii). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(ii), except with CO(ii).2 replaced by (ii). Analogously to case (i), $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(ii).1 by D.1(iii); CO(ii).2 since $(y_{t},x_{t})$ satisfies (ii); and CO(ii).4 via 2.1.
Regarding CO(ii).3, we first note that similarly to (LABEL:alphabetaplus),
@@
so that $\tilde{\alpha}=Q\alpha$ and $\tilde{\beta}^{\pm}=(P^{\pm})^{\mathsf{T}}\beta^{\pm}$, and we may take $\tilde{\alpha}_{\perp}=(Q^{-1})^{\mathsf{T}}\alpha_{\perp}$. With respect to $\tilde{\beta}_{\perp}^{\pm}=\tilde{\beta}_{\perp}(\pm 1)$, for the purposes of verifying CO(ii).3, we need to choose this such that its first row is equal to $e_{q,1}^{\mathsf{T}}$, and its final columns are regime-invariant, as per (LABEL:betaperpfn) (see 4.24.2). This may be engineered by defining
@@
where $\tilde{\beta}_{x,\perp}=\Phi_{0,xx}\beta_{x,\perp}$. Since $\tilde{\Gamma}(1;y)=Q\Gamma(1;y)P(y)$ by D.1(iii), we have
@@
Hence 4.2 applies to $(\tilde{y}_{t},\tilde{x}_{t})$. Regarding the conclusions of that theorem, we first note that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$, $\operatorname{sgn}{\cal Y}_{0}=\operatorname{sgn}{\cal\tilde{Y}}_{0}$,
@@
$Z(\lambda)=P[\tilde{Y}(\lambda)]\tilde{Z}(\lambda)$ $=P[\tilde{Y}(\lambda)]P_{\tilde{\beta}_{\perp}}[\tilde{Y}(\lambda)]\tilde{U}_{0}(\lambda)=P_{\beta_{\perp}}[Y(\lambda)]U_{0}(\lambda).$
@@
Since $\tilde{y}_{t}\sim I^{\ast}(1)$, it follows that $\tilde{y}_{t}^{\pm}\sim I^{\ast}(1)$, and hence $\Delta\tilde{y}_{t}^{\pm}\sim I^{\ast}(0)$. That $\Delta z_{t}\sim I^{\ast}(0)$ then follows as in (LABEL:Diffi0). Deduce that the conclusions of the theorem hold also for $(y_{t},x_{t})$.
(iii). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(iii), except with CO(iii).2 replaced by (ii). Since $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(iii).1 by D.1(iii), it follows from 4.3 that $\tilde{\boldsymbol{\chi}}_{t}\coloneqq(\tilde{\chi}_{t}^{\mathsf{T}},\ldots,\tilde{\chi}_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$ is ${\cal Q}$-geometrically ergodic.
Recalling the definitions in (LABEL:Pydef) above, we may rewrite (LABEL:canon-vars) as
@@
where $\bar{\phi}_{0,xy}(y)\coloneqq-\Phi_{0,xx}^{-1}\phi_{0,xy}(y)\bar{\phi}_{0,yy}^{-1}(y)$. Further, we have by D.1(iii) that
@@
Hence by (LABEL:tildeagain) and (LABEL:tildeabx)
@@
so that the l.h.s. is a Lipschitz continuous function of $\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}$, as claimed. Further, we have from (LABEL:ztoztildeagain) that
@@
Deterministics in cases (i) and (ii). Now suppose that $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$ satisfies DET rather than DET. It follows from (LABEL:projtilde) that $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}\tilde{c}=0$; thus DET also holds for the derived canonical process $\tilde{z}_{t}=(\tilde{y}_{t},\tilde{x}_{t}^{\mathsf{T}})^{\mathsf{T}}$. We then have from the proof of 4.4 (under DGP) that the conclusions of part (ii) of Theorems 4.1 and 4.2 apply to the detrended canonical processes
@@
where $\tilde{y}_{t}^{d}=\tilde{y}_{t}$ because $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}c=0$. Using $P(y)$ as defined in (LABEL:Pydef), and noting that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$, we may write
@@
where $=_{(1)}$ holds because only the first column of $P(y)$ depends on $y$, and $e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c=0$. Hence
@@
where $\tilde{y}_{t}=\tilde{y}_{t}^{d}$. Because the mapping between the detrended processes in (LABEL:zdtotilde) is identical to that for the original processes in (LABEL:ztotilde), the same arguments as given in cases (i) and (ii) immediately above will now transpose the conclusions of 4.4 from $\tilde{z}_{t}^{d}$ to $z_{t}^{d}$.
@@
We first note the following corollary to A.3. We say a VAR is stationary if all its autoregressive roots lie outside the unit circle (i.e. irrespective of whether the series is given a stationary initialisation).
[…]
3
@@
Suppose that the assumptions of A.3 hold, and $\{z_{t}\}$ is generated according to
@@
The claim follows from part (ii) of A.3, and premultiplying the companion form $\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\Pi}\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}$ by $\boldsymbol{\beta}^{\mathsf{T}}$, where $\boldsymbol{z}_{t}=(z_{t}^{\mathsf{T}},z_{t-1}^{\mathsf{T}},\ldots,z_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, $\boldsymbol{c}=(c^{\mathsf{T}},0_{p(k-1)}^{\mathsf{T}})^{\mathsf{T}}$ and $\boldsymbol{u}_{t}=(u_{t}^{\mathsf{T}},0_{p(k-1)}^{\mathsf{T}})^{\mathsf{T}}$. ∎
Proof of 4.14.1.
Rather than working with the matrices $\boldsymbol{F}_{0}$ and $\boldsymbol{F}_{1}$ directly, it is easier if we consider the autoregressive systems that they correspond to, as given in (LABEL:coclas:sm), recognising that those systems are dynamically stable if and only if the eigenvalues of their companion form matrices lie strictly inside the unit circle (e.g. Lütkepohl, 2007, Sec. 2.1.1).
[…]
@@
By A.3, the eigenvalues of $I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}$ lie inside the unit circle; hence so too do those of $\boldsymbol{F}_{0}$. (Indeed, it may be verified that the only nonzero eigenvalues of $\boldsymbol{F}_{0}$ are those of $I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}$.)
We next consider $\boldsymbol{F}_{1}$, which is operative if $\delta_{s}=1$ for all $s\in\{t-k,\ldots,t-1\}$. Then $y_{s}=\text{$\overline{y}$}_{s}=y_{s}^{-}<0$ for all such $s$, and (LABEL:coclas:sm) becomes
@@
and equilibrium errors $\xi_{t}^{-}\coloneqq\beta^{-\mathsf{T}}z_{t}=[\begin{smallmatrix}\beta^{+\mathsf{T}}z_{t}\\ y_{t}\end{smallmatrix}]=[\begin{smallmatrix}\xi_{t}^{+}\\ y_{t}\end{smallmatrix}]$. It follows from E.1 that
@@
evolves according to a stationary VAR. Since $(\boldsymbol{\xi}_{t}^{+},y_{t})$ is simply a reordering of $\boldsymbol{\xi}_{t}^{-}$, it follows that the system described by (LABEL:delta1case), i.e. by $\boldsymbol{F}_{1}$, must also be stationary. Hence $\boldsymbol{F}_{1}$ must have all its eigenvalues strictly inside the unit circle. ∎
Proof of 4.14.1.
Suppose $k=1$. In this case, $\pi^{-}-\pi^{+}=\phi_{1}^{-}-\phi_{1}^{+}$, and so by 4.14.1, the matrix
@@
By A.3, $I_{r}+\beta^{+\mathsf{T}}\alpha^{+}$ has all its eigenvalues inside the unit circle, and so $\det[I_{r}-(I_{r}+\beta^{+\mathsf{T}}\alpha^{+})]>0$. Hence $\kappa_{1}<0$.
@@
Appendix F Detailed calculations for 4.2
Here we provide some further details of the calculations underlying 4.2. For convenience, we reproduce the model in CKSVAR form as
@@
and thus $\Pi^{\pm}$ and $\beta^{\pm}=(0,-1,1)^{\mathsf{T}}$ are regime-invariant. $\Pi^{x}$ is obtained from the final $p-1=2$ columns of $\Pi^{+}$ (or $\Pi^{-}$), and so
@@
In view of (LABEL:PipmABH), we can construct orthocomplement matrices as
@@
where it will be observed that the $\beta_{\perp}^{\pm}$ indeed have the form required by (LABEL:betaperpfn). Since $\Gamma_{i}=0$ for all $i\geq 1$ in this first-order model, we have $\Gamma^{\pm}(1)=\Phi_{0}^{\pm}$, and thus from computing
@@
The final condition to verify is CO(ii).2 – or rather (ii), because the CSKVAR (LABEL:ABH-appendix) is in structural form. We must therefore compute $\tilde{\Pi}^{\pm}$ for the associated canonical form, the parameters can be obtained via the mapping (LABEL:canon-vars)–(LABEL:Q-canon). From (LABEL:PhipmABH) we have (noting that here $y_{t}=\bar{\pi}_{t}$ and $x_{t}=(\Delta y_{t},g_{t})^{\mathsf{T}}$),

### v4

@@ Appendix D Transposition to the structural form
Here we verify the claims made in part (i) of each of Remarks 4.14.3, and complete the proof of Theorem 4.4. We suppose that DGP holds, so that $(y_{t},x_{t})$ are generated by the CKSVAR
@@
By Proposition 2.1, if $(\tilde{y}_{t},\tilde{x}_{t})$ are constructed from $(y_{t},x_{t})$ via (2.13), then $(\tilde{y}_{t},\tilde{x}_{t})$ follow a canonical CKSVAR, which we denote here as
@@
We use tildes to distinguish the series and parameters of the canonical form (D.1), from the original CKSVAR (D.2) from which they were derived; the mapping between these is given in (2.13)–(2.15).
Our approach here is as follows. We show that if DGP, CVAR, DET (or DET) and a suitably modified form of CO(i) (resp. CO(ii), CO(iii)) hold for $(y_{t},x_{t})$, then DGP, CVAR, DET (or DET) and CO(i) (resp. CO(ii), CO(iii)) also hold for $(\tilde{y}_{t},\tilde{x}_{t})$. Since ERR (or ERR) trivially carry over from $u_{t}$ to $\tilde{u}_{t}$, this permits Theorems 4.14.3 to be applied to $(\tilde{y}_{t},\tilde{x}_{t})$: their implications for $(y_{t},x_{t})$ are then derived by inverting the mapping (2.13)–(2.15).
@@
By Proposition 2.1, a CKSVAR satisfying DGP has a canonical form, which relates to the original model via (2.13)–(2.14). Recall the definition of $(P^{\pm})^{-1}$ given in (D.3). Since
@@
as may be verified using the partitioned inverse formula, (2.14) may be equivalently stated as
@@
and thus (i) and (ii) hold. Since $\Pi^{\pm}=-\Phi^{\pm}(1)$ and $\tilde{\Pi}^{\pm}=-\tilde{\Phi}^{\pm}(1)$, (iii) follows immediately upon taking $\lambda=1$ in (D.5), and the fact that
@@
(i). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(i), except with CO(i).2 replaced by (ii). We must verify that $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i). $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i).1 by Proposition D.1(iii); CO(i).2 since $(y_{t},x_{t})$ satisfies (ii); and CO(i).4 via Proposition 2.1.
It remains therefore to verify CO(i).3. We note that as a further consequence of Proposition D.1(iii),
@@
Hence, using Proposition D.1(iii) again,
@@
It follows that Theorem 4.1 applies to $(\tilde{y}_{t},\tilde{x}_{t})$. Regarding the conclusions of that theorem, note that in this case $\tilde{{\cal Y}}_{0}$ and ${\cal Y}_{0}$ must be non-negative. Hence $\tilde{{\cal Z}}_{0}=(P^{+})^{-1}{\cal Z}_{0}$ by (2.13), and so by Proposition 2.1
@@
where $_{(1)}$ follows by (D.8), and $_{(2)}$ by (D.9) and $e_{1}^{\mathsf{T}}(P^{+})^{-1}=\bar{\phi}_{0,yy}^{+}e_{1}^{\mathsf{T}}$. It follows immediately that $\beta^{+\mathsf{T}}z_{t}$ and $y_{t}^{-}$ are $I^{\ast}(0)$. Since $\tilde{y}_{t}^{-}\sim I^{\ast}(0)$, it follows that $\Delta\tilde{y}_{t}^{-}\sim I^{\ast}(0)$, and therefore so too is $\Delta\tilde{y}_{t}^{+}=\Delta\tilde{y}_{t}-\Delta\tilde{y}_{t}^{-}$. Hence
@@
(ii). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(ii), except with CO(ii).2 replaced by (ii). Analogously to case (i), $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(ii).1 by Proposition D.1(iii); CO(ii).2 since $(y_{t},x_{t})$ satisfies (ii); and CO(ii).4 via Proposition 2.1.
Regarding CO(ii).3, we first note that similarly to (D.6),
@@
so that $\tilde{\alpha}=Q\alpha$ and $\tilde{\beta}^{\pm}=(P^{\pm})^{\mathsf{T}}\beta^{\pm}$, and we may take $\tilde{\alpha}_{\perp}=(Q^{-1})^{\mathsf{T}}\alpha_{\perp}$. With respect to $\tilde{\beta}_{\perp}^{\pm}=\tilde{\beta}_{\perp}(\pm 1)$, for the purposes of verifying CO(ii).3, we need to choose this such that its first row is equal to $e_{q,1}^{\mathsf{T}}$, and its final columns are regime-invariant, as per (4.8) (see Remark 4.24.2). This may be engineered by defining
@@
where $\tilde{\beta}_{x,\perp}=\Phi_{0,xx}\beta_{x,\perp}$. Since $\tilde{\Gamma}(1;y)=Q\Gamma(1;y)P(y)$ by Proposition D.1(iii), we have
@@
Hence Theorem 4.2 applies to $(\tilde{y}_{t},\tilde{x}_{t})$. Regarding the conclusions of that theorem, we first note that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$, $\operatorname{sgn}{\cal Y}_{0}=\operatorname{sgn}{\cal\tilde{Y}}_{0}$,
@@
$Z(\lambda)=P[\tilde{Y}(\lambda)]\tilde{Z}(\lambda)$ $=P[\tilde{Y}(\lambda)]P_{\tilde{\beta}_{\perp}}[\tilde{Y}(\lambda)]\tilde{U}_{0}(\lambda)=P_{\beta_{\perp}}[Y(\lambda)]U_{0}(\lambda).$ (D.13)
@@
Since $\tilde{y}_{t}\sim I^{\ast}(1)$, it follows that $\tilde{y}_{t}^{\pm}\sim I^{\ast}(1)$, and hence $\Delta\tilde{y}_{t}^{\pm}\sim I^{\ast}(0)$. That $\Delta z_{t}\sim I^{\ast}(0)$ then follows as in (D.11). Deduce that the conclusions of the theorem hold also for $(y_{t},x_{t})$.
(iii). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(iii), except with CO(iii).2 replaced by (ii). Since $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(iii).1 by Proposition D.1(iii), it follows from Theorem 4.3 that $\tilde{\boldsymbol{\chi}}_{t}\coloneqq(\tilde{\chi}_{t}^{\mathsf{T}},\ldots,\tilde{\chi}_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$ is $\cal Q$}-geometrically ergodic.
Recalling the definitions in (D.12) above, we may rewrite (2.13) as
@@
where $\bar{\phi}_{0,xy}(y)\coloneqq-\Phi_{0,xx}^{-1}\phi_{0,xy}(y)\bar{\phi}_{0,yy}^{-1}(y)$. Further, we have by Proposition D.1(iii) that
@@
Hence by (D.14) and (D.16)
@@
so that the l.h.s. is a Lipschitz continuous function of $\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}$, as claimed. Further, we have from (D.15) that
@@
Deterministics in cases (i) and (ii). Now suppose that $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$ satisfies DET rather than DET. It follows from (D.7) that $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}\tilde{c}=0$; thus DET also holds for the derived canonical process $\tilde{z}_{t}=(\tilde{y}_{t},\tilde{x}_{t}^{\mathsf{T}})^{\mathsf{T}}$. We then have from the proof of Theorem 4.4 (under DGP) that the conclusions of part (ii) of Theorems 4.1 and 4.2 apply to the detrended canonical processes
@@
where $\tilde{y}_{t}^{d}=\tilde{y}_{t}$ because $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}c=0$. Using $P(y)$ as defined in (D.12), and noting that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$, we may write
@@
where $_{(1)}$ holds because only the first column of $P(y)$ depends on $y$, and $e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c=0$. Hence
@@
where $\tilde{y}_{t}=\tilde{y}_{t}^{d}$. Because the mapping between the detrended processes in (D.18) is identical to that for the original processes in (D.17), the same arguments as given in cases (i) and (ii) immediately above will now transpose the conclusions of Theorem 4.4 from $\tilde{z}_{t}^{d}$ to $z_{t}^{d}$.
@@
We first note the following corollary to Lemma A.3. We say a VAR is stationary if all its autoregressive roots lie outside the unit circle (i.e. irrespective of whether the series is given a stationary initialisation).
@@
Suppose that the assumptions of Lemma A.3 hold, and $\{z_{t}\}$ is generated according to
@@
The claim follows from part (ii) of Lemma A.3, and premultiplying the companion form $\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\Pi}\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}$ by $\boldsymbol{\beta}^{\mathsf{T}}$, where $\boldsymbol{z}_{t}=(z_{t}^{\mathsf{T}},z_{t-1}^{\mathsf{T}},\ldots,z_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, $\boldsymbol{c}=(c^{\mathsf{T}},0_{p(k-1)}^{\mathsf{T}})^{\mathsf{T}}$ and $\boldsymbol{u}_{t}=(u_{t}^{\mathsf{T}},0_{p(k-1)}^{\mathsf{T}})^{\mathsf{T}}$. ∎
Proof of Remark 4.14.1.
Rather than working with the matrices $\boldsymbol{F}_{0}$ and $\boldsymbol{F}_{1}$ directly, it is easier if we consider the autoregressive systems that they correspond to, as given in (B.7), recognising that those systems are dynamically stable if and only if the eigenvalues of their companion form matrices lie strictly inside the unit circle (e.g. Lütkepohl, 2007, Sec. 2.1.1).
Let $t\in\mathbb{N}$ be given. We first consider $\boldsymbol{F}_{0}$: this is operative if $\delta_{s}=0$ for all $s\in\{t-k,\ldots,t-1\}$. In this case, $\text{$\overline{y}$}_{s}=y_{s}^{-}=0$ for all such $s$, and (B.7) reduces to
@@
By Lemma A.3, the eigenvalues of $I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}$ lie inside the unit circle; hence so too do those of $\boldsymbol{F}_{0}$. (Indeed, it may be verified that the only nonzero eigenvalues of $\boldsymbol{F}_{0}$ are those of $I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}$.)
We next consider $\boldsymbol{F}_{1}$, which is operative if $\delta_{s}=1$ for all $s\in\{t-k,\ldots,t-1\}$. Then $y_{s}=\text{$\overline{y}$}_{s}=y_{s}^{-}<0$ for all such $s$, and (B.7) becomes
@@
and equilibrium errors $\xi_{t}^{-}\coloneqq\beta^{-\mathsf{T}}z_{t}=[\begin{smallmatrix}\beta^{+\mathsf{T}}z_{t}\\ y_{t}\end{smallmatrix}]=[\begin{smallmatrix}\xi_{t}^{+}\\ y_{t}\end{smallmatrix}]$. It follows from Lemma E.1 that
@@
evolves according to a stationary VAR. Since $(\boldsymbol{\xi}_{t}^{+},y_{t})$ is simply a reordering of $\boldsymbol{\xi}_{t}^{-}$, it follows that the system described by (E.1), i.e. by $\boldsymbol{F}_{1}$, must also be stationary. Hence $\boldsymbol{F}_{1}$ must have all its eigenvalues strictly inside the unit circle. ∎
Proof of Remark 4.14.1.
Suppose $k=1$. In this case, $\pi^{-}-\pi^{+}=\phi_{1}^{-}-\phi_{1}^{+}$, and so by Remark 4.14.1, the matrix
@@
By Lemma A.3, $I_{r}+\beta^{+\mathsf{T}}\alpha^{+}$ has all its eigenvalues inside the unit circle, and so $\det[I_{r}-(I_{r}+\beta^{+\mathsf{T}}\alpha^{+})]>0$. Hence $\kappa_{1}<0$.
@@
Appendix F Detailed calculations for Example 4.2
Here we provide some further details of the calculations underlying Example 4.2. For convenience, we reproduce the system (4.12) in CKSVAR form as
@@
and thus $\Pi^{\pm}$ and $\beta^{\pm}=(0,-1,1)^{\mathsf{T}}$ are regime-invariant. $\Pi^{x}$ is obtained from the final p1=2p-1=2 columns of $\Pi^{+}$ (or $\Pi^{-}$), and so
@@
In view of (F.3), we can construct orthocomplement matrices as
@@
where it will be observed that the $\beta_{\perp}^{\pm}$ indeed have the form required by (4.8). Since $\Gamma_{i}=0$ for all $i\geq 1$ in this first-order model, we have $\Gamma^{\pm}(1)=\Phi_{0}^{\pm}$, and thus from computing
@@
The final condition to verify is CO(ii).2 – or rather (ii), because the CSKVAR (F.1) is in structural form. We must therefore compute $\tilde{\Pi}^{\pm}$ for the associated canonical form, the parameters can be obtained via the mapping (2.13)–(2.15). From (F.2) we have (noting that here $y_{t}=\bar{\pi}_{t}$ and $x_{t}=(\Delta y_{t},g_{t})^{\mathsf{T}}$),